1) Entity N follows the revenue recognition principle.   It services a vehicle on September 30. The customer picks up the vehicle on October 1 and mails the payment to Entity N on October 7. Entity N receives the check in the mail on October 9.   When should Entity N show that the revenue was recognized?

a) september 30th

b) october 7th

c) october 1st

d) october 9th

2. Entity M signed a six-month note payable in the amount of $60,000 on October 1. The note requires interest at an annual rate of 7%. The amount of interest expense Entity M must recognize at December 31 is:

a) $1,050

b) $350

c) $1,200

c) $4,200

3) Which of the following accounting concepts justifies the following practice: Only things that can be expressed in money are included in the accounting records?

A) Full disclosure

B) Consistency

C) Materiality

D) Monetary unit

4) Entity C's total liabilities increased by $75,000 and its stockholders’ equity decreased by $15,000. Entity C's total assets must change by what amount and in what direction during that same period?

A) $60,000 decrease

B) $75,000 increase

C) $60,000 increase

D) $90,000 increase
Answer #1
1
Under revenue recognition principle, revenue is recognized when services are performed.
Entity N Performed the services on September 30.
Entity N should recognize revenue on september 30th
Option A is correct
2
Interest expense at December 31 = 60000*7%*3/12= $1,050
Option A is correct
3
Under the Monetary unit assumption,only things that can be expressed in money are included in the accounting records.
Option D is correct
4
Assets = Liabilities + Stockholders’ equity
Net change in Assets = 75000-15000 = $60,000 increase
Option C is correct